# Unveiling the optimization process of Physics Informed Neural Networks: How accurate and competitive can PINNs be?

### v1

@@ Appendix C Error analysis
$E_{u}^{(p)}=\frac{\lVert u-u_{\mathrm{an}}\rVert_{p}}{\lVert u_{\mathrm{an}}\rVert_{p}},$ (73)
where .$\lVert.\rVert_{p}$ indicates the norm of order $p$. If not, we provide $\epsilon_{\mathrm{NN}}$ as a global estimation of the PINN approximation error, together with an illustrative plot of $\epsilon_{\mathrm{FD}}$ as a function of the resolution $N_{0}$.
Exp. $\left(\tau_{k},\phi_{k}\right)$ $\mathcal{J}$ $E^{(1)}_{\mathcal{P}}$ $E^{(2)}_{\mathcal{P}}$ $E^{(1)}_{B_{r}}$ $E^{(2)}_{B_{r}}$ $E^{(1)}_{B_{\theta}}$ $E^{(2)}_{B_{\theta}}$
$1$ $(1,1)$ $1.69\times 10^{-8}$ $2.34\times 10^{-6}$ $2.09\times 10^{-6}$ $1.16\times 10^{-5}$ $1.29\times 10^{-5}$ $6.36\times 10^{-6}$ $6.34\times 10^{-6}$
$2$ $(1,1)$ $1.40\times 10^{-8}$ $1.86\times 10^{-6}$ $1.81\times 10^{-6}$ $8.38\times 10^{-6}$ $8.59\times 10^{-6}$ $6.13\times 10^{-6}$ $6.57\times 10^{-6}$
$3$ $(1,1)$ $6.23\times 10^{-8}$ $3.45\times 10^{-6}$ $3.38\times 10^{-6}$ $2.88\times 10^{-5}$ $4.09\times 10^{-5}$ $1.15\times 10^{-5}$ $1.15\times 10^{-5}$
$4$ $(1,1)$ $4.75\times 10^{-8}$ $5.45\times 10^{-6}$ $5.05\times 10^{-6}$ $2.30\times 10^{-5}$ $2.74\times 10^{-5}$ $1.22\times 10^{-5}$ $1.36\times 10^{-5}$
$5$ $(1,1)$ $1.41\times 10^{-8}$ $1.55\times 10^{-6}$ $1.46\times 10^{-6}$ $1.20\times 10^{-5}$ $1.55\times 10^{-5}$ $5.06\times 10^{-6}$ $5.17\times 10^{-6}$
Mean $(1,1)$ $3.10\times 10^{-8}$ $2.92\times 10^{-6}$ $2.76\times 10^{-6}$ $1.68\times 10^{-5}$ $2.11\times 10^{-5}$ $8.25\times 10^{-6}$ $8.63\times 10^{-6}$
$1$ $\left(\tau_{k}^{(1)},1\right)$ $2.33\times 10^{-10}$ $2.72\times 10^{-7}$ $2.56\times 10^{-7}$ $2.17\times 10^{-6}$ $3.24\times 10^{-6}$ $5.65\times 10^{-7}$ $3.86\times 10^{-7}$
$2$ $\left(\tau_{k}^{(1)},1\right)$ $1.07\times 10^{-10}$ $1.65\times 10^{-7}$ $1.50\times 10^{-7}$ $1.33\times 10^{-6}$ $1.91\times 10^{-6}$ $4.82\times 10^{-7}$ $5.07\times 10^{-7}$
$3$ $\left(\tau_{k}^{(1)},1\right)$ $4.48\times 10^{-11}$ $1.55\times 10^{-7}$ $1.49\times 10^{-7}$ $9.11\times 10^{-7}$ $1.36\times 10^{-6}$ $3.35\times 10^{-7}$ $3.58\times 10^{-7}$
$4$ $\left(\tau_{k}^{(1)},1\right)$ $1.68\times 10^{-10}$ $1.91\times 10^{-7}$ $1.98\times 10^{-7}$ $1.81\times 10^{-6}$ $2.82\times 10^{-6}$ $4.82\times 10^{-7}$ $4.74\times 10^{-7}$
$5$ $\left(\tau_{k}^{(1)},1\right)$ $1.03\times 10^{-10}$ $2.01\times 10^{-7}$ $1.97\times 10^{-7}$ $1.40\times 10^{-6}$ $2.08\times 10^{-6}$ $4.14\times 10^{-7}$ $3.86\times 10^{-7}$
Mean $\left(\tau_{k}^{(1)},1\right)$ $1.32\times 10^{-10}$ $1.97\times 10^{-7}$ $1.90\times 10^{-7}$ $1.53\times 10^{-6}$ $2.28\times 10^{-6}$ $4.56\times 10^{-7}$ $4.50\times 10^{-7}$
$1$ $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $4.48\times 10^{-11}$ $1.36\times 10^{-7}$ $1.32\times 10^{-7}$ $8.76\times 10^{-7}$ $1.23\times 10^{-6}$ $2.88\times 10^{-7}$ $2.86\times 10^{-7}$
$2$ $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $1.88\times 10^{-11}$ $6.59\times 10^{-8}$ $6.31\times 10^{-8}$ $4.12\times 10^{-7}$ $5.34\times 10^{-7}$ $1.46\times 10^{-7}$ $1.42\times 10^{-7}$
$3$ $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $4.42\times 10^{-11}$ $9.81\times 10^{-8}$ $1.06\times 10^{-7}$ $8.97\times 10^{-7}$ $1.38\times 10^{-6}$ $3.01\times 10^{-7}$ $3.48\times 10^{-7}$
$4$ $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $2.96\times 10^{-11}$ $1.20\times 10^{-7}$ $1.08\times 10^{-7}$ $6.86\times 10^{-7}$ $1.02\times 10^{-6}$ $2.58\times 10^{-7}$ $2.75\times 10^{-7}$
$5$ $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $2.46\times 10^{-11}$ $1.22\times 10^{-7}$ $1.13\times 10^{-7}$ $7.94\times 10^{-7}$ $1.23\times 10^{-6}$ $1.68\times 10^{-7}$ $1.86\times 10^{-7}$
Mean $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $3.24\times 10^{-11}$ $1.08\times 10^{-7}$ $1.04\times 10^{-7}$ $7.33\times 10^{-7}$ $1.07\times 10^{-6}$ $2.32\times 10^{-7}$ $2.48\times 10^{-7}$
Table 4: Final loss values and error norms associated with $\mathcal{P}$ and with the magnetic field components $(B_{r},B_{\theta})$, obtained with the standard BFGS $(1,1)$, the self-scaled BFGS $(\tau_{k}^{(1)},0)$ and the self-scaled Broyden updates $(\tau_{k}^{(2)},\phi_{k}^{(1)})$, for the current-free dipole-quadrupole case.
Exp. Loss $\mathcal{J}$ $E^{(1)}_{\mathcal{P}}$ $E^{(2)}_{\mathcal{P}}$ $E^{(1)}_{B_{r}}$ $E^{(2)}_{B_{r}}$ $E^{(1)}_{B_{\theta}}$ $E^{(2)}_{B_{\theta}}$
$1$ $\mathcal{J}$ $1.69\times 10^{-8}$ $2.34\times 10^{-6}$ $2.09\times 10^{-6}$ $1.16\times 10^{-5}$ $1.29\times 10^{-5}$ $6.36\times 10^{-6}$ $6.34\times 10^{-6}$
$2$ $\mathcal{J}$ $1.40\times 10^{-8}$ $1.86\times 10^{-6}$ $1.81\times 10^{-6}$ $8.38\times 10^{-6}$ $8.59\times 10^{-6}$ $6.13\times 10^{-6}$ $6.57\times 10^{-6}$
$3$ $\mathcal{J}$ $6.23\times 10^{-8}$ $3.45\times 10^{-6}$ $3.38\times 10^{-6}$ $2.88\times 10^{-5}$ $4.09\times 10^{-5}$ $1.15\times 10^{-5}$ $1.15\times 10^{-5}$
$4$ $\mathcal{J}$ $4.75\times 10^{-8}$ $5.45\times 10^{-6}$ $5.05\times 10^{-6}$ $2.30\times 10^{-5}$ $2.74\times 10^{-5}$ $1.22\times 10^{-5}$ $1.36\times 10^{-5}$
$5$ $\mathcal{J}$ $1.41\times 10^{-8}$ $1.55\times 10^{-6}$ $1.46\times 10^{-6}$ $1.20\times 10^{-5}$ $1.55\times 10^{-5}$ $5.06\times 10^{-6}$ $5.17\times 10^{-6}$
Mean $\mathcal{J}$ $3.10\times 10^{-8}$ $2.92\times 10^{-6}$ $2.76\times 10^{-6}$ $1.68\times 10^{-5}$ $2.11\times 10^{-5}$ $8.25\times 10^{-6}$ $8.63\times 10^{-6}$
$1$ $\mathcal{J}_{1/2}$ $5.79\times 10^{-10}$ $4.21\times 10^{-7}$ $4.12\times 10^{-7}$ $3.49\times 10^{-6}$ $5.49\times 10^{-6}$ $1.18\times 10^{-6}$ $1.22\times 10^{-6}$
$2$ $\mathcal{J}_{1/2}$ $3.26\times 10^{-10}$ $2.64\times 10^{-7}$ $2.65\times 10^{-7}$ $1.48\times 10^{-6}$ $1.57\times 10^{-6}$ $1.18\times 10^{-6}$ $1.31\times 10^{-6}$
$3$ $\mathcal{J}_{1/2}$ $8.33\times 10^{-10}$ $4.84\times 10^{-7}$ $4.79\times 10^{-7}$ $3.48\times 10^{-6}$ $4.89\times 10^{-6}$ $1.40\times 10^{-6}$ $1.31\times 10^{-6}$
$4$ $\mathcal{J}_{1/2}$ $9.87\times 10^{-10}$ $4.45\times 10^{-7}$ $4.18\times 10^{-7}$ $4.00\times 10^{-6}$ $5.72\times 10^{-6}$ $1.39\times 10^{-6}$ $1.40\times 10^{-6}$
$5$ $\mathcal{J}_{1/2}$ $2.40\times 10^{-10}$ $2.35\times 10^{-7}$ $2.41\times 10^{-7}$ $1.84\times 10^{-6}$ $2.44\times 10^{-6}$ $5.95\times 10^{-7}$ $5.79\times 10^{-7}$
Mean $\mathcal{J}_{1/2}$ $5.93\times 10^{-10}$ $3.70\times 10^{-7}$ $3.63\times 10^{-7}$ $2.86\times 10^{-6}$ $4.02\times 10^{-6}$ $1.15\times 10^{-6}$ $1.17\times 10^{-6}$
$1$ $\mathcal{J}_{\mathrm{log}}$ $9.19\times 10^{-11}$ $1.09\times 10^{-7}$ $1.03\times 10^{-7}$ $1.13\times 10^{-6}$ $1.66\times 10^{-6}$ $3.93\times 10^{-7}$ $3.99\times 10^{-7}$
$2$ $\mathcal{J}_{\mathrm{log}}$ $9.50\times 10^{-11}$ $1.52\times 10^{-7}$ $1.44\times 10^{-7}$ $1.09\times 10^{-6}$ $1.38\times 10^{-6}$ $4.25\times 10^{-7}$ $4.14\times 10^{-7}$
$3$ $\mathcal{J}_{\mathrm{log}}$ $3.42\times 10^{-10}$ $3.05\times 10^{-7}$ $2.92\times 10^{-7}$ $2.33\times 10^{-6}$ $3.16\times 10^{-6}$ $8.17\times 10^{-7}$ $7.69\times 10^{-7}$
$4$ $\mathcal{J}_{\mathrm{log}}$ $6.33\times 10^{-10}$ $3.66\times 10^{-7}$ $3.64\times 10^{-7}$ $2.83\times 10^{-6}$ $3.85\times 10^{-6}$ $1.11\times 10^{-7}$ $1.17\times 10^{-6}$
$5$ $\mathcal{J}_{\mathrm{log}}$ $2.62\times 10^{-10}$ $2.73\times 10^{-7}$ $2.52\times 10^{-7}$ $1.63\times 10^{-6}$ $2.01\times 10^{-6}$ $9.87\times 10^{-7}$ $1.01\times 10^{-6}$
Mean $\mathcal{J}_{\mathrm{log}}$ $2.85\times 10^{-10}$ $2.41\times 10^{-7}$ $2.31\times 10^{-7}$ $1.80\times 10^{-6}$ $2.41\times 10^{-6}$ $7.47\times 10^{-7}$ $7.55\times 10^{-7}$
Table 5: Final loss values and error norms associated with $\mathcal{P}$ and with the magnetic field components $(B_{r},B_{\theta})$ obtained with the standard BFGS $(1,1)$ in conjunction with the loss function modifications defined in (27), (28) for the current-free dipole-quadrupole case.
Exp. $\left(\tau_{k},\phi_{k}\right)$ $\epsilon_{\mathrm{NN}}$
1 $(1,1)$ $1.57\times 10^{-3}$
2 $(1,1)$ $1.53\times 10^{-3}$
3 $(1,1)$ $1.98\times 10^{-3}$
4 $(1,1)$ $1.93\times 10^{-3}$
5 $(1,1)$ $1.77\times 10^{-3}$
Mean $(1,1)$ $1.76\times 10^{-3}$
1 $\left(\tau_{k}^{(1)},1\right)$ $3.45\times 10^{-5}$
2 $\left(\tau_{k}^{(1)},1\right)$ $5.42\times 10^{-5}$
3 $\left(\tau_{k}^{(1)},1\right)$ $5.39\times 10^{-5}$
4 $\left(\tau_{k}^{(1)},1\right)$ $2.47\times 10^{-5}$
5 $\left(\tau_{k}^{(1)},1\right)$ $3.15\times 10^{-5}$
Mean $\left(\tau_{k}^{(1)},1\right)$ $3.98\times 10^{-5}$
1 $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $1.45\times 10^{-5}$
2 $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $1.84\times 10^{-5}$
3 $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $1.44\times 10^{-5}$
4 $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $1.29\times 10^{-5}$
5 $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $1.51\times 10^{-5}$
Mean $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $1.51\times 10^{-5}$
Exp. Loss $\epsilon_{\mathrm{NN}}$
1 $\mathcal{J}$ $1.57\times 10^{-3}$
2 $\mathcal{J}$ $1.53\times 10^{-3}$
3 $\mathcal{J}$ $1.98\times 10^{-3}$
4 $\mathcal{J}$ $1.93\times 10^{-3}$
5 $\mathcal{J}$ $1.77\times 10^{-3}$
Mean $\mathcal{J}$ $1.76\times 10^{-3}$
1 $\mathcal{J}_{1/2}$ $4.77\times 10^{-4}$
2 $\mathcal{J}_{1/2}$ $3.10\times 10^{-4}$
3 $\mathcal{J}_{1/2}$ $4.06\times 10^{-4}$
4 $\mathcal{J}_{1/2}$ $2.79\times 10^{-4}$
5 $\mathcal{J}_{1/2}$ $4.14\times 10^{-4}$
Mean $\mathcal{J}_{1/2}$ $3.98\times 10^{-4}$
1 $\mathcal{J}_{\mathrm{log}}$ $6.68\times 10^{-5}$
2 $\mathcal{J}_{\mathrm{log}}$ $5.40\times 10^{-5}$
3 $\mathcal{J}_{\mathrm{log}}$ $9.29\times 10^{-5}$
4 $\mathcal{J}_{\mathrm{log}}$ $1.33\times 10^{-4}$
5 $\mathcal{J}_{\mathrm{log}}$ $2.97\times 10^{-5}$
Mean $\mathcal{J}_{\mathrm{log}}$ $6.93\times 10^{-5}$
Table 6: Values of $\epsilon_{\mathrm{NN}}$ obtained for the eight multipoles force-free solution. In 6(a) they are calculated by considering the BFGS modifications together with the standard MSE loss $\mathcal{J}$, whereas in 6(b) by considering the loss function modifications in conjunction with the standard BFGS algorithm.
Exp. $\left(\tau_{k},\phi_{k}\right)$ $\mathcal{J}$ $E_{\phi}^{(2)}$
1 $\left(1,1\right)$ $\mathcal{J}$ $1.71\times 10^{-6}$
2 $\left(1,1\right)$ $\mathcal{J}$ $2.55\times 10^{-6}$
3 $\left(1,1\right)$ $\mathcal{J}$ $2.06\times 10^{-6}$
Mean $\left(1,1\right)$ $\mathcal{J}$ $2.11\times 10^{-6}$
1 $\left(1,1\right)$ $\mathcal{J}_{\mathrm{log}}$ $1.69\times 10^{-8}$
2 $\left(1,1\right)$ $\mathcal{J}_{\mathrm{log}}$ $7.81\times 10^{-8}$
3 $\left(1,1\right)$ $\mathcal{J}_{\mathrm{log}}$ $7.60\times 10^{-8}$
Mean $\left(1,1\right)$ $\mathcal{J_{\mathrm{log}}}$ $5.70\times 10^{-8}$
1 $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $\mathcal{J}$ $1.91\times 10^{-8}$
2 $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $\mathcal{J}$ $3.52\times 10^{-8}$
3 $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $\mathcal{J}$ $4.75\times 10^{-8}$
Mean $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $\mathcal{J}$ $3.39\times 10^{-8}$
Table 7: Values of $\epsilon_{\mathrm{NN}}$ obtained with the standard BFGS $(1,1)$, the self-scaled BFGS $(\tau_{k}^{(1)},0)$ and the self-scaled Broyden updates $(\tau_{k}^{(2)},\phi_{k}^{(1)})$, for the Non-linear Poisson solution.
Figure 14: $\epsilon_{\mathrm{FD}}$ as a function of the resolution $N_{0}$ for the equation (38). A very similar trend is also obtained for the equation (39)
Exp. $\left(\tau_{k},\phi_{k}\right)$ Loss $\epsilon_{\mathrm{NN}}$ (38) $\epsilon_{\mathrm{NN}}$ (39)
1 $(1,1)$ $\mathcal{J}$ $3.89\times 10^{-4}$ $3.57\times 10^{-4}$
2 $(1,1)$ $\mathcal{J}$ $4.54\times 10^{-4}$ $4.19\times 10^{-4}$
3 $(1,1)$ $\mathcal{J}$ $4.32\times 10^{-4}$ $4.15\times 10^{-4}$
Mean $(1,1)$ $\mathcal{J}$ $\bm{4.25\times 10^{-4}}$ $\bm{3.97\times 10^{-4}}$
1 $(1,1)$ $\mathcal{J}_{\mathrm{log}}$ $3.84\times 10^{-5}$ $3.38\times 10^{-5}$
2 $(1,1)$ $\mathcal{J}_{\mathrm{log}}$ $3.65\times 10^{-5}$ $3.25\times 10^{-5}$
3 $(1,1)$ $\mathcal{J}_{\mathrm{log}}$ $4.34\times 10^{-5}$ $3.69\times 10^{-5}$
Mean $(1,1)$ $\mathcal{J}_{\mathrm{log}}$ $\bm{3.94\times 10^{-5}}$ $\bm{3.34\times 10^{-5}}$
1 $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $\mathcal{J}$ $2.66\times 10^{-5}$ $2.17\times 10^{-5}$
2 $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $\mathcal{J}$ $1.54\times 10^{-5}$ $1.32\times 10^{-5}$
3 $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $\mathcal{J}$ $2.42\times 10^{-5}$ $2.41\times 10^{-5}$
Mean $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $\mathcal{J}$ $\bm{2.20\times 10^{-5}}$ $\bm{1.96\times 10^{-5}}$
Table 8: Values of $\epsilon_{\mathrm{NN}}$ obtained for the non-linear Schrödinger equation, considering the standard BFGS algorithm with $\mathcal{J}$, the standard BFGS algorithm with $\mathcal{J}_{\mathrm{log}}$, and the self-scaled Broyden algorithm with $\mathcal{J}$.
Exp. $\left(\tau_{k},\phi_{k}\right)$ $\mathcal{J}$ $E_{u}^{(2)}$
1 $\left(1,1\right)$ $\mathcal{J}$ $1.97\times 10^{-4}$
2 $\left(1,1\right)$ $\mathcal{J}$ $1.43\times 10^{-4}$
3 $\left(1,1\right)$ $\mathcal{J}$ $9.08\times 10^{-4}$
Mean $\left(1,1\right)$ $\mathcal{J}$ $4.16\times 10^{-4}$
1 $\left(1,1\right)$ $\mathcal{J}_{\mathrm{log}}$ $1.43\times 10^{-5}$
2 $\left(1,1\right)$ $\mathcal{J}_{\mathrm{log}}$ $1.44\times 10^{-5}$
3 $\left(1,1\right)$ $\mathcal{J}_{\mathrm{log}}$ $7,84\times 10^{-6}$
Mean $\left(1,1\right)$ $\mathcal{J_{\mathrm{log}}}$ $1.22\times 10^{-5}$
1 $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $\mathcal{J}$ $7.26\times 10^{-6}$
2 $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $\mathcal{J}$ $7.00\times 10^{-6}$
3 $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $\mathcal{J}$ $5,96\times 10^{-6}$
Mean $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $\mathcal{J}$ $6.74\times 10^{-6}$
Table 9: Error norms of the PINN approximation for the KdV two soliton solution, obtained with the standard BFGS and the MSE loss, with the standard BFGS and $\mathcal{J}_{\mathrm{log}}$ defined at (28), and with the self-scaled Broyden updates $(\tau_{k}^{(2)},\phi_{k}^{(1)})$ in conjunction with the MSE loss.
Exp. $\left(\tau_{k},\phi_{k}\right)$ Loss $E_{u}^{(2)}$ $E_{v}^{(2)}$ $E_{w}^{(2)}$ $E_{p}^{(2)}$
1 $(1,1)$ $\mathcal{J}$ $1.23\times 10^{-5}$ $1.56\times 10^{-5}$ $1.44\times 10^{-5}$ $1.08\times 10^{-4}$
2 $(1,1)$ $\mathcal{J}$ $9.67\times 10^{-6}$ $1.27\times 10^{-5}$ $1.34\times 10^{-5}$ $9.69\times 10^{-5}$
3 $(1,1)$ $\mathcal{J}$ $1.69\times 10^{-5}$ $1.93\times 10^{-5}$ $2.10\times 10^{-5}$ $1.76\times 10^{-4}$
Mean $\left(1,1\right)$ $\mathcal{J}$ $1.30\times 10^{-5}$ $1.59\times 10^{-5}$ $1.63\times 10^{-5}$ $1.27\times 10^{-4}$
1 $(1,1)$ $\mathcal{J}_{\mathrm{log}}$ $1.13\times 10^{-6}$ $1.35\times 10^{-6}$ $1.27\times 10^{-5}$ $1.27\times 10^{-5}$
2 $(1,1)$ $\mathcal{J}_{\mathrm{log}}$ $1.29\times 10^{-6}$ $1.49\times 10^{-6}$ $1.50\times 10^{-6}$ $1.44\times 10^{-5}$
3 $(1,1)$ $\mathcal{J}_{\mathrm{log}}$ $1.38\times 10^{-6}$ $1.25\times 10^{-6}$ $1.22\times 10^{-6}$ $1.32\times 10^{-5}$
Mean $\left(1,1\right)$ $\mathcal{J}_{\mathrm{log}}$ $1.27\times 10^{-6}$ $1.36\times 10^{-6}$ $1.33\times 10^{-6}$ $1.34\times 10^{-5}$
1 $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $\mathcal{J}$ $6.65\times 10^{-7}$ $6.75\times 10^{-7}$ $7.98\times 10^{-7}$ $7.17\times 10^{-6}$
2 $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $\mathcal{J}$ $8.53\times 10^{-7}$ $6.76\times 10^{-7}$ $7.73\times 10^{-7}$ $7.50\times 10^{-6}$
3 $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $\mathcal{J}$ $6.55\times 10^{-7}$ $7.69\times 10^{-7}$ $7.64\times 10^{-7}$ $6.65\times 10^{-6}$
Mean $\left(\tau_{k}^{(2)},\phi_{k}^{(1)}\right)$ $\mathcal{J}$ $7.24\times 10^{-7}$ $7.07\times 10^{-7}$ $7.78\times 10^{-7}$ $7.06\times 10^{-6}$
Table 10: Error norms for the three velocity components and the pressure obtained with the standard BFGS and the MSE loss, with the standard BFGS and $\mathcal{J}_{\mathrm{log}}$ defined at (28), and with the self-scaled Broyden updates $(\tau_{k}^{(2)},\phi_{k}^{(1)})$ in conjunction with the MSE loss.

### v2

@@ Appendix C Error analysis
$\epsilon_{\mathrm{FD}}=\frac{1}{N_{0}^{d/2}}\left\lVert\triangle_{\mathrm{GS}}^{\mathrm{FD}}\mathcal{P}+\mathcal{T}\left(\frac{d\mathcal{T}}{d\mathcal{P}}\right)\right\rVert_{2}=\frac{1}{N_{0}^{d/2}}\sqrt{\sum_{ij}\Big{|}\triangle_{\mathrm{GS}}^{\mathrm{FD}}\mathcal{P}_{ij}+\mathcal{T}_{ij}\left(\frac{d\mathcal{T}}{d\mathcal{P}}\right)_{ij}\Big{|}^{2}},$ (90)
where $d$ is the dimension of the problem (2D in this example). If $\mathcal{P}_{ij}$ were the exact solution of the PDE, the error $\epsilon_{\mathrm{FD}}$ would decrease with increasing resolution as $\sim N_{0}^{-p}$, with $p$ being the order of the finite difference approximation. In reality, $\mathcal{P}_{ij}$ is only an approximate solution with an intrinsic error $\epsilon_{\mathrm{NN}}$ depending on the (unknown) accuracy of the PINN. A reasonable way to measure $\epsilon_{\mathrm{NN}}$ is to examine the convergence of the discretized PDE residuals with resolution. $\epsilon_{\mathrm{FD}}$ will follow the $\sim N_{0}^{-p}$ power law only up to the point where the PINN approximation error becomes dominant. Beyond this point, $\epsilon_{\mathrm{FD}}$ will level off and remain roughly constant with increasing resolution, approximately equal to $\epsilon_{\mathrm{NN}}$. This behavior is clearly seen, for example, in Fig. 4. If the PINN approximation is inaccurate, this will happen at a relatively low resolution. To quantify the PINN approximation error using this approach, we define
$\epsilon_{\mathrm{NN}}=\min\left\{\epsilon_{\mathrm{FD}}\right\}.$ (91)
It is worth noting that, with this strategy of using finite differences to validate the PINNs solution, the intrinsic errors due to the approximation of the derivatives are always measurable and can be reduced as much as needed (typically is a power law with increasing resolution). One can simply use a higher-order formula involving more points or even use quadruple precision if needed. This is precisely the main advantage of this validation method: it is simple and its accuracy is always under control, unlike the usual comparison with direct results from simulations which are also subject to (sometimes uncertain) errors, and there are situations where it is unclear whether the difference between the PINNs results and the numerical results are due to the PINN inaccuracy or to some other limitation in the code used as benchmark.